# Segment then Splat: A Unified Approach for 3D Open-Vocabulary Segmentation based on Gaussian Splatting

### v1

@@ 3.5 CLIP Embedding Association
3.6 Open-Vocabulary Query
Given an input text prompt, we perform open vocabulary query following the below strategy:
@@ 4 Experiments
4.1 Dataset
To assess the segmentation performance of our proposed method, we conduct experiments on two static scene datasets (i.e., 3DOVS dataset [15] and LERF_OVS dataset [9]) and two dynamic scene datasets (i.e., HyperNeRF dataset [20] and Plenoptic Video dataset [13]).
4.2 Implementation Details
The new object detection stride $\Delta t$ in the robust object tracking module is set to 10. Following the original 3DGS implementation, we set $\lambda_{r}$ in $\mathcal{L}_{\text{render}}$ to 0.2. For $\lambda_{d}$, which balances geometric distance and appearance distance in Eq. 4, we use a value of 0.5. During each iteration, we sample one object per granularity for 3DOVS to compute $\mathcal{L}_{\text{obj}}$ and three objects per granularity for all the remaining datasets. The mIoU threshold for partial mask filtering is set to 70%. For the 3DOVS dataset, given its smaller scale, we train for 20,000 iterations (5,000 for stage1, 5,000 for stage2, and the remaining for stage3), and introducing partial mask filtering in the last 5,000 iterations. For larger-scale datasets (i.e. LERF_OVS, HyperNeRF, and Plenoptic), training extends to 40,000 iterations (5,000 for stage1, 5,000 for stage2, and the remaining for stage3), with partial mask filtering applied during the last 10,000 iterations.
4.3 Baselines and Evaluation Metrics
Baselines. We categorize the baselines into two groups based on their querying strategies: 2D pixel-based segmentation and 3D-based segmentation. For static scenes, we use LangSplat [21] and LEGaussians [21] as the 2D pixel-based baselines. For the 3D baselines, we choose OpenGaussian[32], and we also adapt LangSplat and LEGaussians for 3D segmentation to evaluate their performance. In dynamic scenes, we adopt a zero-shot image segmentation model CLIP-LSeg [12] as the 2D pixel-based baseline and DGD [11] as the 3D-based approach.
Evaluation Metrics. We adopt mean intersection over union (mIoU) as evaluation metric for open-vocabulary segmentation task. For training efficiency, since all baselines and our proposed method share a similar preprocessing pipeline, we only focus on comparing the optimization time. All the time in the tables are measured in minutes. One thing to be mentioned is that, due to the nature of 3D segmentation and the evaluation used for open-vocabulary segmentation, there is a slight reduction in our mIoU score. This is because 3D segmentation directly retrieves the Gaussians associated with an object, revealing occluded parts that are not visible in some ground truth masks, as shown in Fig. 5.
@@ 4.4 Open-Vocabulary Query
Results on 3DOVS dataset. The quantitative results on the 3DOVS dataset are presented in Tab. 1. Our method outperforms all baseline approaches. The qualitative comparison is shown in Fig. 4. Notably, 2D pixel-based methods tend to produce relatively ambiguous boundaries, whereas our approach, leveraging the “segment then splat” strategy, achieves significantly clearer object boundaries. For 3D segmentation methods, OpenGaussian requires a predefined number of objects $K$ for clustering. Since 3DOVS contains fewer objects compared to LERF_OVS, we manually reduced its preset $K$. However, even after this adjustment, OpenGaussian still only retrieves parts of certain objects. Although LangSplat and LEGaussians are modified to segment Gaussians instead of pixels, their performance remains suboptimal. This is because the language embedding of each Gaussian is not directly supervised and may encode multiple semantics from different objects, leading to inaccurate segmentation results. Additionally, since our method follows a single-pass reconstruction process and the scene scale is relatively small, our training time is significantly shorter compared to the baseline methods.
Table 1: Quantitative segmentation results on static scenes.
@@
Results on LERF_OVS dataset. The quantitative results on the LERF_OVS dataset are also presented in Tab. 1 and the qualitative results are shown in Fig. 4. Similar to the 3DOVS dataset, 2D pixel-based methods produce less precise object boundaries, while our method demonstrates significantly improved results. For 3D-based methods, OpenGaussian performs much better on LERF_OVS compared to 3DOVS. However, since OpenGaussian requires a predefined number of clusters, some objects may be incorrectly grouped together (e.g., the rubber duck and the Rubik’s cube). Moreover, due to its “splat then segmentation” strategy, its object boundaries remain less accurate than those produced by our Segment then Splat method.
Table 2: Quantitative segmentation results on dynamic scenes.
HyperNeRF Planoptic
@@
Results on HyperNeRF dataset. The quantitative results and qualitative results on HyperNeRF dataset are presented in Tab. 2 and Fig. 6 respectively. Since our method explicitly enforces Gaussian-object correspondence, it can be directly applied to dynamic scenes, achieving good segmentation performance without the Gaussian-object misalignment issue encountered by previous approaches. In contrast, methods such as DGD, which relies on learning a language field supervised via 2D feature maps, suffer from misalignment, as a single Gaussian might represent multiple distinct objects across different time steps. As illustrated in Fig. 6, this misalignment results in retrieving irrelevant Gaussians. Moreover, because DGD does not directly supervise the language embeddings of each Gaussian, Gaussians located far apart may share similar embeddings, further deteriorating segmentation quality. In addition, our method achieves nearly a ten-fold improvement in optimization speed compared to DGD, as learning a dynamic language field is computationally intensive. We omit training time results for LSeg, as it is a zero-shot segmentation method requiring no additional optimization.
Results on Plenoptic Video dataset. Quantitative results and qualitative results are shown in Tab. 2 and Fig. 6. Similar to the observations on the HyperNeRF dataset, many irrelevant Gaussians are retrieved due to object-Gaussian misalignment issue and the “splat then segment” strategy. Besides, both LSeg and DGD fail when retrieving small objects (e.g. bunny painting).
Figure 6: Qualitative comparison on dynamic scenes. As our method enforce object-Gaussian correspondence, it can be directly applied to dynamic scenes and achieve good performance, whereas DGD and LSeg tend to include irrelevant content and are not able to retrieve small objects (e.g. bunny painting).
@@ 4.5 Ablation Study
Number of Objects to Supervise. In this ablation study, we examine the relationship between the number of supervised objects in each granularity per iteration and the segmentation performance. We pick two static scenes from LERF_OVS dataset and two dynamic scenes from HyperNeRF dataset. The results are presented in Tab. 3. As expected, increasing the number of supervised objects per iteration generally enhances segmentation performance. However, beyond a certain threshold, the performance gain becomes marginal. For scenes with fewer objects (e.g. chickchicken and split-cookie), performance quickly converges after a certain number of supervised objects. In contrast, more complex scenes containing many objects (e.g. ramen and waldo_kitchen) continue to show performance improvements, though at a diminished rate. Additionally, training time increases proportionally with the number of supervised objects. To balance training time and segmentation performance, we choose three objects for LERF_OVS, Plenoptic and HyperNeRF in our experiments.
Table 3: Ablation study on the number of supervised objects per iteration at each granularity level. As the number of supervised objects increases, we observe an improvement in segmentation performance, but also an increase in training time.
Static Scene Dynamic Scene
ramen waldo_kitchen chickchicken split-cookie
[…]
1 26 51.09 44 33.97 146 73.11 180 77.76
3 34 54.38 48 40.71 157 74.89 202 80.30
5 42 55.61 54 40.83 172 75.29 253 80.47
7 53 56.03 63 40.77 185 75.21 258 81.52
9 62 56.48 68 41.59 201 75.23 309 81.74
Partial Mask Filtering. In this ablation study, we investigate the impact of the partial mask filtering strategy on both segmentation performance and reconstruction quality. As shown in Tab. 4, scenes with a larger number of objects (e.g., ramen and waldo_kitchen) tend to have more occluded views, leading to a more significant performance improvement when applying partial mask filtering. In contrast, scenes with fewer objects (e.g., chickchicken and split-cookie) exhibit a smaller performance gain. Overall, the results validate the effectiveness of our proposed partial mask filtering strategy in improving segmentation accuracy, particularly in complex scenes with high object density.
Table 4: Ablation study on the partial mask filtering strategy, w/o MF stands for without partial mask filtering.
Static Scene Dynamic Scene
ramen waldo_kitchen chickchicken split-cookie
[…]
w/o MF 42.19 23.11 31.94 30.43 72.65 29.32 80.23 32.76
ours 54.38 24.54 40.71 31.33 74.89 29.62 80.30 33.04
@@ 5 Conclusion
We propose Segment then Splat, a unified framework for 3D open-vocabulary segmentation based on Gaussian Splatting. We reverse the long-established “segmentation after reconstruction” approach to form a “segmentation then reconstruction” pipeline. By enforcing object-Gaussian correspondence through the entire process, Segment then Splat yields a better object boundary and can be directly applied to dynamic scenes. Finally, open-vocabulary querying is enabled by assigning each object-specific Gaussian group with a unified CLIP embedding. Extensive experiments on various static and dynamic scene datasets demonstrate the superior performance of our method.

### v2

@@ 3.5 CLIP Embedding Association
Open-vocabulary segmentation. Given an input text prompt, we perform open vocabulary query following the below strategy:
@@ 4 Experiments
4.1 Setups
Baselines. We categorize the baselines into two groups based on their querying strategies: 2D pixel-based segmentation and 3D-based segmentation. For static scenes, we use LangSplat [10], LEGaussians [11] and Gaussian Grouping [36] as the 2D pixel-based baselines. For the 3D baselines, we choose OpenGaussian[13], and we also adapt LangSplat and LEGaussians for 3D segmentation by selecting Gaussians instead of pixels to evaluate their performance. In dynamic scenes, we adopt a zero-shot image segmentation model CLIP-LSeg [29] as the 2D pixel-based baseline and DGD [28] as the 3D-based approach.
Datasets & Evaluation Metrics. To assess the segmentation performance of our proposed method, we conduct experiments on two static scene datasets (i.e., 3DOVS dataset [37] and LERF_OVS dataset [9]) and two dynamic scene datasets (i.e., HyperNeRF dataset [38] and Neu3D dataset [39]). We use mean intersection over union (mIoU) for open-vocabulary segmentation and report optimization time (in minutes) for training efficiency.
Implementation Details. The new object detection stride $\Delta t$ in the robust object tracking is set to 10. Following the original 3D Gaussian Splatting, we set $\lambda_{r}$ in $\mathcal{L}_{\text{render}}$ to 0.2. For geometric-appearances distance, we set $\lambda_{d}$ to 0.5. In each iteration, we sample 1 object per granularity for 3DOVS to compute $\mathcal{L}_{\text{obj}}$ and 3 objects per granularity for all the remaining datasets. The mIoU threshold for partial mask filtering is set to 30%. We train the smaller-scale 3DOVS data for 20K iterations, and larger-scale datasets (i.e., LERF_OVS, HyperNeRF, and Neu3D) for 40K iterations. All experiments are conducted using a RTX A6000 GPU.
@@ 4.2 Open-Vocabulary Query
2D V.S. 3D segmentation. One thing to be mentioned is that, due to the nature of 3D segmentation and the evaluation used for open-vocabulary segmentation, there is a slight reduction in our mIoU score. This is because 3D segmentation directly retrieves the Gaussians associated with an object, revealing occluded parts that are not visible in some ground truth masks, as shown in Fig.˜5.
Table 1: Quantitative segmentation results for static (a) and dynamic (b) scenes.
(a) Static scenes
@@
(b) Dynamic scenes
HyperNeRF Neu3D
@@
Results on 3DOVS dataset. The quantitative results on the 3DOVS are presented in Tab.˜1 (a). Our method outperforms all baseline approaches. The qualitative comparison is shown in Fig.˜4. Notably, 2D pixel-based methods tend to produce relatively ambiguous boundaries, whereas our approach, leveraging the Segment then Splat strategy, achieves significantly clearer object boundaries. For 3D segmentation methods, OpenGaussian requires a predefined number of objects $K$ for clustering. Since 3DOVS contains fewer objects compared to LERF_OVS, we manually reduced its preset $K$. However, even after this adjustment, OpenGaussian still only retrieves parts of certain objects. Although LangSplat and LEGaussians are modified to segment Gaussians instead of pixels, their performance remains suboptimal. This is because each Gaussian’s language embedding lacks direct supervision and may encode multiple object semantics, leading to inaccurate segmentation. Additionally, since our method follows a single-pass reconstruction process and the scene scale is relatively small, our training time is significantly shorter compared to the baselines.
Results on LERF_OVS dataset. The quantitative results on the LERF_OVS dataset are also presented in Tab.˜1 (a) and the qualitative results are shown in Fig.˜4. Similar to the 3DOVS dataset, 2D pixel-based methods produce less precise object boundaries, while our method demonstrates significantly improved results. For 3D-based methods, OpenGaussian performs much better on LERF_OVS compared to 3DOVS. However, since OpenGaussian requires a predefined number of clusters, some objects may be incorrectly grouped together (e.g., the rubber duck and the Rubik’s cube). Moreover, due to its “splat then segmentation” strategy, its object boundaries remain less accurate than those produced by our method.
Results on HyperNeRF dataset. The quantitative results and qualitative results on HyperNeRF dataset are presented in Tab.˜1 (b) and Fig.˜6 respectively. Since our method explicitly enforces Gaussian-object correspondence, it can be directly applied to dynamic scenes, achieving good segmentation performance without the Gaussian-object misalignment issue encountered by previous approaches. In contrast, methods such as DGD, which relies on learning a language field supervised via 2D feature maps, suffer from misalignment, as a single Gaussian might represent multiple distinct objects across different time steps. As illustrated in Fig.˜6, this misalignment results in retrieving irrelevant Gaussians. Moreover, because DGD does not directly supervise the language embeddings of each Gaussian, Gaussians located far apart may share similar embeddings, further deteriorating segmentation quality.
Table 2: Ablation studies: (Top) Number of supervised objects per iteration. (Middle) Partial mask filtering. (Bottom) Tracking module ablation.
(a) Number of supervised objects per iteration
[…]
(b) Partial mask filtering strategy
Static Scene Dynamic Scene ramen waldo_kitchen chickchicken split-cookie Method $uparrow mIoU\uparrow PSNR\uparrow$ $uparrow mIoU\uparrow PSNR\uparrow$ w/o MF 42.19 23.11 31.94 30.43 72.65 29.32 80.23 32.76 Ours 54.38 24.54 40.71 31.33 74.89 29.62 80.30 33.04
(c) Tracking module ablation
[…]
In addition, our method achieves nearly a ten-fold improvement in optimization speed compared to DGD, as learning a dynamic language field is computationally intensive. We omit training time results for LSeg, as it is a zero-shot method requiring no additional optimization.
Results on Neu3D dataset.
Quantitative results and qualitative results are shown in Tab.˜1 and Fig.˜6. Similar to the observations on the HyperNeRF dataset, many irrelevant Gaussians are retrieved due to object-Gaussian misalignment issue and the “splat then segment” strategy. Besides, both LSeg and DGD fail when retrieving relatively small objects (e.g., bunny painting).
@@ 4.3 Ablation Study
Figure 6: Qualitative comparison on dynamic scenes. As our method enforce object-Gaussian correspondence, it applies directly to dynamic scenes and performs well, whereas DGD and LSeg tend to include irrelevant content and are not able to retrieve small objects (e.g., bunny painting).
Number of Supervised Objects. In this ablation study, we examine the relationship between the number of supervised objects in each granularity per iteration and the segmentation performance. We pick two static scenes from LERF_OVS and two dynamic scenes from HyperNeRF dataset. The results are presented in Tab.˜2 (a). As expected, increasing the number of supervised objects per iteration generally enhances segmentation performance. However, beyond a certain threshold, the performance gain becomes marginal. For scenes with fewer objects (e.g., chickchicken and split-cookie), performance quickly converges after a certain number of supervised objects. In contrast, more complex scenes containing many objects (e.g., ramen and waldo_kitchen) continue to show performance improvements, though at a diminished rate. Additionally, training time increases proportionally with the number of supervised objects. To balance training time and performance, we choose three objects for LERF_OVS, Neu3D and HyperNeRF in our experiments.
Partial Mask Filtering. In this ablation study, we investigate the impact of the partial mask filtering strategy on segmentation performance as well as reconstruction quality. As shown in Tab.˜2 (b), scenes with a larger number of objects (e.g., ramen, waldo_kitchen) tend to have more occluded views, leading to more performance gain when applying partial mask filtering. In contrast, scenes with fewer objects (e.g., chickchicken, split-cookie) exhibit a smaller performance gain. Overall, the results validate the effectiveness of the partial mask filtering strategy, particularly in complex scenes with high object density.
Robust Object Tracking. We conduct an ablation study on each component of our robust object tracking module, as shown in Tab.˜2 (c). The evaluation is performed on three static scenes from LERF_OVS and one dynamic scene from HyperNeRF. Leveraging ground-truth labels, we adopt two metrics: Object Recall Rate (ORR), defined as
$\text{ORR}=\frac{1}{k}\sum_{i=1}^{k}\frac{\text{number of tracked objects}}{\text{number of GT objects}},$ (12)
where k is the number of ground-truth frames, measures the percentage of objects successfully tracked throughout the sequence; and the number of duplicate trackings (Dup), which quantifies how often the same object is redundantly tracked as multiple instances. Results show that the new object detection module increases ORR by capturing newly appeared objects, albeit at the cost of more duplicate instances. In contrast, the multi-track and lost-track handling modules effectively reduce Dup, as described in Sec.˜3.2. The slight drop in ORR observed in the Figurines scene when adding the multi-track module is caused by an isolated tracking failure at a fine-grained level in a single frame. However, this minor failure does not affect the final reconstruction, as sufficient information is retained from other views.
@@ 5 Conclusion
We propose Segment then Splat, a unified framework for 3D open-vocabulary segmentation based on Gaussian Splatting. We reverse the long-established “segment after reconstruct” approach to form a “segment then reconstruct” pipeline. By maintaining consistent object–Gaussian correspondence throughout the process, Segment then Splat effectively eliminates both geometric and semantic ambiguities, resulting in more precise object boundaries and improved segmentation performance. Furthermore, because this correspondence is explicitly established, our method naturally extends to dynamic scenes without concerns of misalignment between objects and Gaussians during dynamic modeling. Extensive experiments across diverse static and dynamic datasets demonstrate the superior performance and robustness of our framework.
6 Limitation and Future Direction
Our method relies on SAM2 for initializing object tracking. In extremely complex scenarios with high-density and visually similar objects, tracking may fail, resulting in suboptimal open-vocabulary segmentation, which is an issue commonly shared by tracking-based approaches (e.g., Gaussian Grouping). This limitation could potentially be addressed by incorporating techniques such as Kalman filtering [40] to improve the temporal stability of SAM2. Another limitation is that our method cannot effectively handle text queries involving relational descriptions across multiple objects, such as “a sheep sitting on the chair in front of the table.” Similar to existing baselines, our method encodes textual embeddings from masked regions or local patches corresponding to individual objects, and therefore lacks explicit modeling of inter-object relationships or contextual understanding. Developing this capability is crucial for advancing open-vocabulary understanding, and we plan to explore it in future work.